OpenStudy (anonymous):

75 cos ^2 x- 34 cos x + 3 =0

OpenStudy (anonymous):

factorise the equation in term of (ax+c)(bx+d)=0

OpenStudy (anonymous):

i dont know how to factor

OpenStudy (anonymous):

if u hav Ax^2 + Bx +C = 0 then, x= -B/2a + sqrt(B^2-4AC)/2A or x= -B/2A - sqrt(B^2-4AC)/2A

OpenStudy (anonymous):

so using the quadratic formula with that problem how would you do that

OpenStudy (anonymous):

u substitute the A,B,C from ur problem into the formulae

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

compare: 75 cos ^2 x- 34 cos x + 3 = Ay^2 + By +C therefore: A=75, B=-34, C= 3 and y=cos x